dy/dx=-cos(xy)+xy sin(xy)/2y-x^2 sin(xy)


1
Expert's answer
2020-10-08T16:28:05-0400

Given differential Equation is "\frac{dy}{dx} = \frac{-cos(xy)+xysin(xy)}{2y-x^2sin(xy)}"


It can be written as, "(cos(xy)-xysin(xy))dx +(2y-x^2sin(xy))dy = 0"

Comparing with "Mdx +Ndy = 0"

For exact differential equation,

"\frac{\partial N}{\partial x} =\frac{\partial M}{\partial y}"

Then

"\frac{\partial N}{\partial x} = -2xsin(xy) -x^2ycos(xy)"


"\frac{\partial M}{\partial y} = -2xsin(xy) -x^2ycos(xy)"


Hence equation is exact differential equation.


Integrating both sides,

"\int (cos(xy)-xysin(xy))dx +\int (2y-x^2sin(xy))dy = C"


"\frac{sin(xy)}{y} - \frac{sin(xy)}{y} + xcos(xy) + y^2 = C"


"xcos(xy) + y^2 = C"